For how long would £500 have to be left in an account paying 4% interest p.a. to give a balance of £600?




What is asked:




What are given:




What is the answer:

1
Expert's answer
2021-11-15T18:30:53-0500

We are looking for the time it would take "\$500"  in an account paying "4\%" interest p.a. to give a balance of "\$600" ,and we are given;


Principal="\$500" rate=0.04, A="\$600"


Simple interest ,A=P(1+rt)


600 =500(1+0.04t)


1.2-1=0.04t


t="\frac{0.2}{0.04}=5"


"\therefore" it would take 5 years for £500 in an account paying 4% interest p.a. to give a balance of £600